Good Morning,
I'm after a bit of help, trying to help my son with his algebra homework but im not to sure on algebra.
Question
1) Find at least on sequence between each pair of numbers, fully describe the term-term rule you have used.
a) 1...9 b) 1...16 c) 5...21 d)5...21 e) 10...20 f) 16...8
I would just like some understanding what needs to be done to find the answer so I can help my son.
Your help would be greatly appreciated.
Many Thanks,
Sarah

Answer by josgarithmetic(39618)

Decide on a rule that will start with the first number and will give the last number. The question does not restrict you from finding numbers between the first and last, but any numbers between the first and last must still conform to the rule.

Items c and d have the same starting and ending numbers. This must mean that you are expected to use a different sequence for each item. Same first and last numbers, but different rules. For simplicity, you might want to choose common difference between terms. Five is an odd number; and 21 is an odd number. Try an even number difference. Notice 21-5=16 so there are just a few differences between terms to choose.

Starting with 5, if you add 8, and then use this as a common difference, then
a list of these terms will be 5+0*8=5; 5+1*8=13; 5+2*8=21.

Starting with 5, if you add 2, then you can use 2 as the common difference and then a list of terms will be 5, 7, 9, 11, 13, 15, 17, 19, 21.